Find the first five nonzero terms in the solution of the given initial value problem. y + xy + 2y = 0, y(0) = 5, y (0) = 3
Answer #1

Here assuming y as a series in x, with the help of given inial values and differential equation; we get the desired solution.

Let y = annn be the series solution, n=0 of the given differential equation. differentiating. y= E nannni then on word ni n=Criven initial value 46.) = 5 and y(o)=3 Now y(0) = 5 = got Gn.o. n=1 9=5 and y () = 3 = Enan.o + 19, na2 9. = 3 Thus we haZ n(n-1) an nn-2 tn nannh- n=2 n=1 + 2 2 anan anan = 0 no I n(n-1) an nh-2 + nannh h=2 h + 2 annno hao Now we can write by reAbus 5 (142) (11) 2 1) + รีก 40 hon1 + 10 hะ *C 0 = 44 % ดsu 9 2.): + 2(142)(+) 90+1 72 + = ร.ศ. ๒-๓ +124 17 + 2 .40 = 0 > 2Put n=1 +2) 9,=0 (1+2) (1+az + (1+2) 693 +3. (3)=0 (:9=3) az = -3/2 Put n=2 (2+2)(2+1) 94 +(2+2) 9,50 (-5)=0 12 94 + 4 (: 4,5